# LLMs in the SOC: An Empirical Study of Human-AI Collaboration in Security Operations Centres

### v1

@@ 5.4 Limitations and Future Work
This study has several limitations. First, all data were collected from a single SOC using a specific LLM deployment. SOCs vary in tools, workflows, and policies, so our findings may not generalise to other operational contexts or different LLMs. Second, we lack objective performance metrics, such as triage speed, false‐positive rates, LLM response accuracy, or overall investigation accuracy. However, it is important to note that the primary goal of this study was not to evaluate performance outcomes, but to explore how analysts interacted with the LLM in practice. High engagement levels suggest analysts found the LLM useful, but we cannot conclude whether these interactions translated into better or faster decisions. Third, there may be a novelty effect, with some analysts engaging with the LLM simply because it was new or conveniently accessible, rather than due to demonstrated utility. As such, it is difficult to distinguish genuine productivity gains from usage driven by curiosity or experimentation without further inquiry.
Future work can build on our findings in several ways. First, replicating this study across multiple SOCs with different toolchains, team sizes, and operating procedures will help assess the generalisability of observed interaction patterns and support the use of objective measures, such as time‐to‐triage. Second, integrating LLM features directly into SOC dashboards could reduce context-switching and offer deeper insight into how interface design shapes analyst-LLM interaction. Finally, longitudinal studies that track new and experienced analysts over time can reveal how trust, adoption, and productivity evolve as the novelty of the LLM wears off. By pursuing these directions, future work can move from descriptive usage patterns and provide more comprehensive insights into the role of LLMs in supporting analyst decision-making and collaboration within operational SOCs.
Ethical Consideration: Ethical approval for this study was granted by the researchers’ host organisation’s Ethics and Privacy Office (Approval #: 025/25). Analysts were informed that their interactions would be anonymously logged to understand how they engaged with the new LLM tools, and consented to this. Informed consent was obtained by the industry partner, and participation was completely voluntary. Prior to data sharing, the industry partner de-identified the data and removed sensitive information. All procedures were reviewed and approved by both the industry partner and the host organisation’s ethics and privacy teams.
@@ 6 Conclusion
This paper explored how SOC analysts interact with LLMs in an operational setting. By analysing thousands of analyst-generated queries, we found that analysts use LLMs as on-demand, task-focused cognitive aids for a variety of tasks, including explaining commands, writing scripts, or improving documentation, rather than as full-time copilots. Future tools should embed LLMs directly into SOC systems to deliver adaptive, context-sensitive assistance. From a research perspective, our work provides a grounded foundation for understanding analyst-LLM collaboration and underscores the need for outcome-based evaluations of their real-world impact on SOC decision-making.

### v2

@@ 5.5 Limitations and Future Work
This study has several limitations. First, all data were collected from a single SOC using a specific LLM deployment. SOCs vary in tools, workflows, and policies, so our findings may not generalise to other operational contexts or different LLMs. Second, we lack objective performance metrics, such as triage speed, false‐positive rates, LLM response accuracy, or overall investigation accuracy. However, it is important to note that the primary goal of this study was not to evaluate performance outcomes but to explore how analysts interacted with the LLM in practice. High engagement levels suggest analysts found the LLM useful, but we cannot conclude whether these interactions translated into better or faster decisions. Third, there may be a novelty effect, with some analysts engaging with the LLM simply because it was new or conveniently accessible, rather than due to demonstrated utility, and some may have avoided it because of the perceived hallucination issues. As such, it is hard to separate genuine productivity gains from curiosity-driven use without further investigation.
Future work can build on our findings in several ways. First, replicating this study across multiple SOCs with different toolchains, team sizes, and operating procedures will help assess the generalisability of observed interaction patterns and support the use of objective measures, such as time‐to‐triage. Second, integrating LLM features directly into SOC dashboards could reduce context-switching and offer deeper insight into how interface design shapes analyst-LLM interaction. Finally, longitudinal studies that track new and experienced analysts over time can reveal how trust, adoption, and productivity evolve as the novelty of the LLM wears off. Future longitudinal studies should include interviews to capture analysts’ experiences and perspectives, integrating these insights into design guidelines for effective LLM-analyst collaboration. By pursuing these directions, future work can move from descriptive usage patterns and provide more comprehensive insights into the role of LLMs in supporting analyst decision-making and collaboration within operational SOCs.
Ethical Considerations: Ethical approval for this study was granted by the researchers’ host organisation’s Ethics and Privacy Office (Approval #: 025/25). Analysts were informed that their interactions would be anonymously logged to understand how they engaged with the new LLM tools, and consented to this. Informed consent was obtained by the industry partner, and participation was completely voluntary. Prior to data sharing, the industry partner de-identified the data and removed sensitive information. All procedures were reviewed and approved by both the industry partner and the host organisation’s ethics and privacy teams.
@@ 6 Conclusion
This paper explored how SOC analysts interact with LLMs in an operational setting. By analysing thousands of analyst-generated queries, we found that analysts use LLMs as on-demand, task-focused cognitive aids for a variety of tasks, including explaining commands, writing scripts, or improving documentation, rather than as full-time copilots. Future tools could embed LLMs directly into SOC systems to deliver adaptive, context-sensitive assistance. From a research perspective, our work provides a grounded foundation for understanding analyst-LLM collaboration and underscores the need for outcome-based evaluations of their real-world impact on SOC decision-making.
@@ 7 Acknowledgments
Appendix A Paper Appendix: Themes and Additional LLM Usage Patterns
LLM Usage Summary: Monthly Patterns. Figure 8 summarises SOC analyst engagement with the LLM system, showing monthly active users (top) and new vs. returning users (bottom), highlighting adoption trends and increasing integration into daily workflows.
Figure 8: Analysts’ engagement over months.
Figure 9: Monthly query volume by task theme.
Monthly Task‐Theme Usage. Figure 9 shows monthly query volumes by Task Theme, highlighting broad growth that suggests LLMs have become multi-role assistants in SOC investigations, reporting, and automation.
TABLE III: Summary of Pattern Theme with Top 3 Codes
Theme Theme Definition Top 3 Code Patterns (Count, %) # Codes (Total)
Functional Understanding Queries aimed at understanding the behaviour, role, or functionality of code, commands, etc what does [x] do (311, 46.7%), what is [x] doing (68, 10.2%), what is this [x] doing (32, 4.8%) 129 (666)
Text Processing, Editing, Summarising Prompts ask the LLM to improve, rephrase, or refine given text make this better [x] (68, 13.8%), reword [x] (25, 5.1%), make it better [x] (21, 4.3%) 214 (494)
Definition, Clarification or Categorisation Requests Questions seek definitions, descriptions, or classifications of a concept, term, or entity what is [x] (186, 41.0%), what is this [x] (38, 8.4%), what are [x] (16, 3.5%) 137 (454)
Explanation, Justification, Causal Reasoning Requests for reasoning or explanations about what something does or why it behaves a certain way. explain this [x] (97, 23.8%), tell me about [x] (26, 6.4%), explain [x] (21, 5.1%) 181 (408)
Content Generation & Expansion Prompts instruct the LLM to create, add to, or expand content such as code or text create [x] (24, 11.4%), can you give me [x] (19, 9.0%), write [x] (12, 5.7%) 105 (211)
LLM Interaction & System Queries Queries and responses directed at the system itself, including gratitude or system-related questions thank you [x] (40, 24.8%), thanks [x] (25, 15.5%), What [x] of gpt are you (7, 4.3%) 58 (161)
Confirmation or Open-ended used to either confirm a known or suspected fact (yes/no, true/false) or to explore a topic is [x] [x] (9, 6.5%), what are your thougths on [x] (3, 2.2%), [x] True or False (3, 2.2%) 114 (138)
How-To, Problem-Solving & Procedural Queries about performing a task, solving a problem, or carrying out a procedure how can i [x] (6, 4.8%), how to [x] (5, 4.0%), in [x] pick [x] (4, 3.2%) 99 (124)
Feasibility & Capability Checking Queries testing whether the LLM can perform a certain action or process a specific input can you [x] (5, 5.5%), can you [x] the following [x] (3, 3.3%), does [x] support [x] (2, 2.2%) 80 (91)
Correction & Error Fixing To identify, explain, or resolve errors in outputs, responses, or code can you correct the [x] (62, 69.7%), correct [x] (5, 5.6%), can you correct [x] (4, 4.5%) 19 (89)
Data Analysis and Interpretation Prompts focused on decoding, interpreting, or drawing insights from data or patterns decode [x] (15, 22.7%), how is [x] (3, 4.5%), are these [x] same (2, 3.0%) 48 (66)
Information Gathering & Retrieval Requests for factual information, examples, or references related to a topic or concept tell me about [x] (2, 3.4%), Is there [x] (2, 3.4%), give me [x] (2, 3.4%) 52 (58)
Explore Options & Alternatives Questions that explore multiple possibilities, compare options, or consider different approaches to a problem, decision, or scenario. Often involves weighing alternatives or identifying the most appropriate course of action. which of the following [x] (2, 3.8%), i need [x] for the [x] (1, 1.9%), which [x] patched [x] (1, 1.9%) 52 (53)
Event & Outcome Analysis Aimed at interpreting security events, determining what occurred, assessing impact, or understanding post-activity implications (e.g., after a login, installation, or process execution) what is the outcome of [x] (6, 15.8%), whats happening here [x] (4, 10.5%), what is the result of [x] (3, 7.9%) 27 (39)
Observation and Analyst Feedback Statements or follow-up queries where analysts reflect on or react to prior LLM responses, sometimes providing their own assessment, confirming correctness, or probing further clarification im trying to [x] could [x] help with that (1, 2.6%), but the previous version of the [x] worked (1, 2.6%), I can’t find any information to verify [x] (1, 2.6%) 38 (38)
TABLE IV: Summary of Task Theme with Top 3 Codes
Theme Theme Definition Top 3 Code Patterns (Count, %) # Codes (Total)
Command Understanding, Analysis & Generation Centred on interpreting, analysing, or generating command-line inputs, including their structure, function, and potential threat relevance Command Understanding/Analysis (861, 90.8%), Command Threat Analysis (25, 2.6%), Command Writing (17, 1.8%) 21 (948)
Text Generation & Editing Involves the generation, rewriting, or refinement of text, including grammar correction Editing/Rewriting Text (130, 19.2%), Grammar Correction (76, 11.2%), Improving Incident Description (31, 4.6%) 116 (676)
Code, Script, and Regex Analysis Involves writing, reading, or interpreting code, scripts, and regular expressions across various programming or automation contexts Code Writing and Understanding (128, 37.5%), Script Understanding (44, 12.9%), Understanding ML concept (24, 7.0%) 41 (341)
Tool / Software / Application / System Understanding Focused on gaining insight into how specific tools, software platforms, or system configurations work Tool Understanding (54, 33.1%), Software Understanding (14, 8.6%), Registry Understanding (5, 3.1%) 62 (163)
LLM Chat: Greeting, Thanking, Version, Memory, Auditing Centred around greetings, expressions of thanks, comments on versioning, or audit-related feedback LLM Interaction (Thanking) (75, 48.1%), LLM Interaction (Greeting) (20, 12.8%), Casual Chat/Conversation (15, 9.6%) 25 (156)
File & Malware Analysis Related to understanding file attributes, locations, and content, often in the context of detecting malicious behaviour File Understanding (94, 61.4%), File Location Analysis (12, 7.8%), File-Related Threat Analysis (6, 3.9%) 30 (153)
Attack, Security & Threat Intelligence Related to understanding attacks, malware types, vulnerabilities, and broader security threats for contextualisation or reporting Attack Understanding (24, 17.0%), Trojan Understanding (7, 5.0%), Attack Type Understanding (6, 4.3%) 75 (141)
Query, Search, & Data Processing Involves the creation, refinement, or execution of search queries, often for use in log analysis, data filtering, or threat hunting Query Writing (72, 60.0%), Answering Multiple Choice Questions (Query) (8, 6.7%), Query Understanding (6, 5.0%) 23 (120)
Network and Traffic Analysis Focused on analysing IP addresses, URLs, and network traffic to understand communication patterns, identify threats, or interpret network behaviour Network Traffic Understanding (18, 19.6%), IP Address Analysis (14, 15.2%), URL Analysis (8, 8.7%) 42 (92)
Access Control & Authentication Concerns remote access protocols, authentication methods, and policy or permission management Policy & Access Management (12, 13.0%), Remote Management Protocol Understanding (10, 10.9%), Authentication Method Understanding (9, 9.8%) 45 (92)
Email Security & Analysis Involving assessment of email rules, phishing detection, and analysis of potentially malicious email behaviours like forwarding, auto-deletion Email Rule Understanding (21, 34.4%), Email Rule Threat Assessment (13, 21.3%), Email Rule Analysis (5, 8.2%) 16 (61)
Alert & Log Analysis Tasks focused on interpreting security alerts, analysing logs and event messages Log Analysis (4, 11.8%), Log Event Analysis (3, 8.8%), Windows Event Management (3, 8.8%) 24 (34)
Others Miscellaneous Finding Movie Word Counts in a Range (5, 14.7%), General Knowledge Related Task (4, 11.8%), Keyword Research (SEO) (2, 5.9%) 21 (34)
Concept Definition To define, clarify, or understand security-related concepts, systems, or identity mechanisms in depth Concept Definition (7, 23.3%), Term Understanding (6, 20.0%), Concept Understanding (4, 13.3%) 15 (30)
Document/Evidence Search Aimed at retrieving or summarising reports, templates, or locating relevant intelligence sources or documentation to support investigative work Identifying Supporting Reports/Papers (13, 61.9%), Obtaining a Report (4, 19.0%), Fact Checking (1, 4.8%) 6 (21)
System Process Understanding & Analysis Involving analysis of system-level behaviours, understanding inter-process communication or OS-level services OS Process Understanding (3, 21.4%), Process Activity Investigation (3, 21.4%), Process & IPC Analysis (2, 14.3%) 8 (14)
Numerical Analysis Involving math-related problem solving, percentage calculations, or interpreting numerically expressed patterns Math Problem (6, 42.9%), Numerical Analysis (3, 21.4%), Math Problem Notation (1, 7.1%) 7 (14)
TABLE V: Summary of Subject Theme with Top 3 Codes
Theme Theme Definition Top 3 Code Patterns (Count, %) # Codes (Total)
command related (specific command, command line interfaces, multiple commands) About command-line syntax, utilities, or individual commands used in operating systems or security-related tasks command (831, 88.9%), command - malicious (9, 1.0%), command line utility (8, 0.9%) 60 (935)
message / paragraph / phrase / sentence / term / article / report / templates / formatting instructions About written content such as phrases, sentences, messages, or templates, including wording, formatting, terminology, and stylistic adjustments message/phrase/paragraph/sentence (277, 53.4%), previous llm response (40, 7.7%), term (21, 4.0%) 105 (519)
coding and scripting related (specific language, function, formula, code snippet, implementations, libraries, packages, development environment, errors, ecoding types, regex) About code, scripts, or regular expressions, including syntax, structure, or functionality across various programming or scripting languages script (41, 14.3%), code (28, 9.8%), coding error (10, 3.5%) 155 (286)
software & tool related (different types, OS, OS-level software, apps, applications, device drivers, development environments, installations, processes, OS-level processes, tools) About specific software tools and applications, including features, setup options, or security-related configurations tool (14, 6.2%), software (10, 4.4%), specific bitlocker configuration - bitlocker policy (9, 4.0%) 150 (225)
file and directory related (including naming conventions, types, file actions, permissions, file behaviours, sharing mechanism, directory, directories related) About file names, file paths, directories, and file system locations relevant to detection or investigation file (file name) (82, 52.9%), file (file name) - specific location/folder (17, 11.0%), directory (3, 1.9%) 51 (155)
LLM/GPT related (including privacy, chat aduditing, version, history) Referencing gpt or large language models, including their versions, capabilities, or role-based configurations llm_gpt (119, 79.9%), gpt version (9, 6.0%), persona description (6, 4.0%) 15 (149)
attack & security related (attack type, indicators, attacker motivations, concepts, solutions, scam, phishing, incidents, APT groups, threat groups, malware, malicious activity) About malicious activity, including different types of malware (e.g., trojans, ransomware), attack techniques (e.g., phishing, brute force), and associated vulnerabilities that enable exploitation malware type (12, 8.7%), attack type (12, 8.7%), network scanning method (4, 2.9%) 94 (138)
MDR & Query related (including edr, edr query, rules, ids, rules, configuration, log related) About queries, filters, or rules related to managed detection and response (MDR) systems or log analysis platforms query (20, 16.5%), sumologic query (9, 7.4%), Sumo logic (6, 5.0%) 70 (121)
alert, investigation, incident, security event related About security alerts, incident details, investigation steps, or descriptions of suspicious or confirmed events incident description (33, 30.8%), alert description (29, 27.1%), investigation description (13, 12.1%) 22 (107)
network communication mechanisms related (including sockets, http, ftp, ports, API, rdp, packet data, protocol, ip address, id address range, cidr) About network-level elements such as IP addresses, port numbers, and remote access protocols like RDP port number (7, 6.7%), ip address (6, 5.7%), regex - ip address (3, 2.9%) 77 (105)
email and mail flow management related (email rules, forwarding, mail flow config, client communication: emails, alerts, incidents, replies) Different types of email rules and their interpretation email (22, 26.8%), email rule (11, 13.4%), email rules (8, 9.8%) 38 (82)
Education and Interview related (training materials, Q&A, interview questions) Content focused on training, self-assessment, and preparation for interviews, such as scenario-based questions scenario based interview questions (15, 20.5%), mitre-related questions and answer (13, 17.8%), different questions and detailed answers (10, 13.7%) 22 (73)
access management related (users, accounts, groups, roles, privileges, active directory, authentication) Topics related to user identities, roles, groups, privileges, authentication methods, and account activity authentication methods (2, 3.1%), authentication method - specific os (1, 1.5%), organisation units - gpo - hosts (1, 1.5%) 64 (65)
general data types (numbers, strings, data, time, hash, hex, base64, table, lists, json) Numbers, strings, data, time, hash, hex, base64, table, lists, JSON decoding data (10, 17.2%), math problem (6, 10.3%), number division (3, 5.2%) 30 (58)
online sources and corroborating evidence or source related (research, papers, reports, online sources) Evidence that could backup LLM or analyst url (9, 30.0%), corroborating evidence or source - research (2, 6.7%), url click event (2, 6.7%) 19 (30)
general terms Miscellaneous or off-topic queries movie word counts (5, 23.8%), world event (4, 19.0%), joke (4, 19.0%) 10 (21)
device and hardware identify related (host, hostnames, domain names, storage device information) Identifiers and information about hardware, hosts, domains, USB devices, or endpoints domains (3, 25.0%), usb details (2, 16.7%), device (1, 8.3%) 9 (12)
organisation and agreement related (companies, competitors, organisational policies and agreements) Organisational context, including companies, competitors, compliance requirements, policies compliance/agreement (2, 22.2%), company competitors (2, 22.2%), internal organisation process (1, 11.1%) 7 (9)